OpenStudy (anonymous):

whats teh simplest form of >9 ÷ 2one-fourth im iving out medals and i have answered the rest please and thank you to those who answer

OpenStudy (anonymous):

\[9\div 2\tfrac{1}{4}=9\div \frac{9}{4}=9\times \frac{4}{9}\]

OpenStudy (anonymous):

i got it just before u answered anyways thanks